OpenStudy (anonymous):

can you help me with this question? Use the Law of Cosines to solve the problem. A ship travels due west for 83 miles. It then travels in a northwest direction for 111 miles and ends up 165 miles from its original position. To the nearest tenth of a degree, how many degrees north of west did it turn when it changed direction? Show your work.

OpenStudy (anonymous):



hartnn (hartnn):

did you try this ? do you know the Law of cosines ?

OpenStudy (anonymous):

no but i read it and don't understand it?

hartnn (hartnn):

here you need angle x , use law of cosines,  \(c^2=a^2+b^2-2ab \cos x\) just plug in a,b,c, values!